OpenStudy (anonymous):

The height of a cell phone tower, AB, is 220 feet.

OpenStudy (anonymous):

What is the distance, in feet, between the tip of the tower A and a point C on the ground? (4 points) 220 sin 46° 220 sec 46° 220/tan46 220/cosec46 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

cos(46) = adj/hyp cos(46) = 220/hyp hyp = 220/cos(46) hyp = 220*sec(46)

jimthompson5910 (jim_thompson5910):

So the distance from A to C is 220*sec(46)